OpenStudy (nannie):

Try this again How woud you multiply & simplify, by factoring? All expressions under radicals are positive. ³√y^13 ³√16y^14

OpenStudy (anonymous):

\[\sqrt[3]{y^{13}}*\sqrt[3]{16y^{14}}=\sqrt[3]{2^4y^{27}}\]note 3 divide 4 1 time with a remainder of 1; 3 divides 27 3 times with no remainder, so may write\[=\sqrt[3]{2^3*2^1(y^9)^3}\]the two bases with the exponents of 3 will come outside the radical, the base with the exponent of 1 will stay inside\[=2y^9*\sqrt[3]{2}\]

OpenStudy (anonymous):

wow you know good math..

OpenStudy (anonymous):

:})

OpenStudy (anonymous):

can you help me with my other question please?

OpenStudy (nannie):

Thanks!!!